OpenStudy (anonymous):

The area of a rectangle is 65 m2 , and the length of the rectangle is 3 m less than twice the width. Find the dimensions of the rectangle.

OpenStudy (dayakar):

let width of a rectangle = x m lenth= twice the width -3 =( 2x-3)m

OpenStudy (dayakar):

area = 65 m^2 given L * W = 65 (2x-3)x =65 2x^2-3x-65=0 2x^2 +10x-13x-65=0 2x(x-5)-13(x-5)=0 (x-5)(2x-13) =0 x-5=0 or 2x-13=0

OpenStudy (dayakar):

x=5 or x=13/2 width x= 5m or x= 6.5m then find length